The rate of motorcycle is 40 mph greater than the rate of a bicycle. The motorcycle travels 150 miles in the same amount of time as the bicycle travels 30 miles. Find the rate of the motocycle.

two people working together can do a job in 3 hrs. How long will it take the slower person to do the same job if one of them is 3 times as fast as the other?

Answer by mangopeeler07(462)

The rate of motorcycle is 40 mph greater than the rate of a bicycle. The motorcycle travels 150 miles in the same amount of time as the bicycle travels 30 miles. Find the rate of the motocycle.

m=rate of motorcycle
b=rate of bike
m=b+40

150/b+40=30/b-----------------------because miles over mph gives you the time, which is equal in this case.

Cross multiply
150b=30b+1200

Subtract 30b from both sides
120b=1200

Divide both sides by 120
b=10

m=b+40
m=10+40
m=50

Answer:
Bike=10mph
Motorcycle=50mph
----------------------------------------------------------------------------
two people working together can do a job in 3 hrs. How long will it take the slower person to do the same job if one of them is 3 times as fast as the other?

together=1/3 (of the job in an hour)
one person=x
other person=3x

x+3x=1/3

Combine like terms
4x=1/3

Divide both sides by 4
x=1/12

1/12 of the job in one hour, so the whole job in 12 hours.

Answer: The slower person can do the job alone in 12 hours.
* The faster person can do the job alone in 4 hours.